whem i subtract 6 from my favorite number, then divide the difference by 3, and finally add 12 to this result, the answer is 20. what is my favorite number?

Answer by Paul(988)

Work the problem backwards.
Add changes to subbtract and divison changes to multiplication.
Difference changes to addition.
20-12=8
8*3=24
24+6=30
Hence, the number is 30.
Paul.